# Context Robust Knowledge Editing for Language Models

### v1

@@ Abstract
Knowledge editing (KE) methods offer an efficient way to modify knowledge in large language models. Current KE evaluations typically assess editing success by considering only the edited knowledge without any preceding contexts. In real-world applications, however, preceding contexts often trigger the retrieval of the original knowledge and undermine the intended edit. To address this issue, we have developed CHED—a benchmark designed to evaluate the context robustness of KE methods. Evaluations on CHED show that they often fail when preceding contexts are present. To mitigate this shortcoming, we introduce CoRE, a KE method designed to strengthen context robustness by minimizing context-sensitive variance in hidden states of the model for edited knowledge. This method not only improves the editing success rate in situations where a preceding context is present but also preserves the overall capabilities of the model. We also provide an in-depth analysis of the differing impacts of preceding contexts when introduced as user utterances versus assistant responses, and we dissect attention-score patterns to assess how specific tokens influence editing success. We release our dataset and code at https://github.com/holi-lab/CoRE.
Context Robust Knowledge Editing for Language Models
@@ 2 Related Work
As a first step toward context robust editing, our CoRE method focuses on the locate-then-edit paradigm due to its capacity for supporting significant edits while maintaining model performance. However, we acknowledge the importance of examining and improving context robustness of editing methods in other paradigms and leave this to future work.
@@ 3.2 Contextual Word Selection (Figure 2-B)
We explored additional criteria to identify words that are closely and uniquely associated with the entities in given fact triplets. For instance, we considered hop words with high cosine similarity to the main entity based on BERT embeddings, capturing semantic closeness. Additionally, we measured the probability that a hop word co-occurs with the entity. Table 2 summarizes the criteria used for hop word selection.
@@
Additionally, we quantitatively assessed the coherence using G-Eval Liu et al. (2023) with GPT-4o-mini. The average coherence score across six types of prefix contexts is 3.4 on a scale from 1 to 5, indicating moderate plausibility. Importantly, if we exclude the $o^{*}$ and $o^{*}_{\!hop}$ types, whose content necessarily diverges from the real world facts, the mean coherence rises to 3.8. Moreover, the mean score does not increase further because we selected primarily low-frequency hop words (as detailed in §3.2), which inherently constrained coherence potential; however, since editing success must remain effective under any preceding context, we deliberately prioritized distractiveness and consider this level of coherence acceptable. A detailed explanation of the G-Eval process and results is provided in Appendix A.7.
@@ 4.1 Preliminaries
Transformer MLP as a Key–Value Associative Memory
MEMIT interprets MLP layers in Transformers as linear associative memories Anderson (1972); Kohonen (1972), where the weights of the projection layer store key–value associations. For example, when a prompt such as “Tim Cook, who works for” is provided as input, the hidden state of the subject’s last token (i.e., “Cook”) encoded by the first MLP layer serves as the key vector $\mathbf{k}$. As $\mathbf{k}$ passes through the second MLP layer $W_{\text{proj}}$, the stored association relevant to the subject is retrieved and embedded into the output value vector $\mathbf{v}$ that contains information about the associated object (e.g., Apple). At subsequent layers, attention mechanisms refine and propagate this recalled knowledge from the value vector, leading the model to generate the token for $o$ Meng et al. (2022a); Geva et al. (2023).
@@
Key–Value Vector Extraction
A key challenge is constructing the key–value pairs that encode the factual edit $(s,r,o)\to(s,r,o^{*})$. $\mathbf{k}$ and $\mathbf{v}$ are derived from a prompt $p$ that includes $s$ and $r$. In MEMIT, various prefix contexts $x_{j}$ are prepended to $p$ to improve contextual generalization. Given $N$ prefix contexts, the key vector is derived as $\mathbf{k}=\frac{1}{N}\sum_{j=1}^{N}k(x_{j}+p),$ where $k(\cdot)$ is obtained by extracting the MLP activation at the last subject token from a chosen layer. We defer the full derivation to Appendix B.1.
@@ 5.2 Experimental Settings
In this paper, we focus on comparing locate-then-edit methods, as they reliably handle a large number of edits, including JEEP Shi et al. (2024), EMMET Gupta et al. (2024b), and PMET Li et al. (2023). We also include FT-M Zhang et al. (2024b) as a representative fine-tuning approach. While we experimented with two representative approaches—a meta-learning method MEND Mitchell et al. (2022a) and a weight-preserved method IKE Zheng et al. (2023a)—both achieved only $\sim\!0\!-\!1\%$ edit success under our stricter generation-based metric, effectively amounting to complete editing failure on 1,000 edits. Consequently, we omit them from Table 4. Details about each method, their results, and hyperparameter settings are provided in Appendix F.7.
@@ 5.3 Main Results
Recent language models are typically trained for dialogues with the user using instruction templates Touvron et al. (2023); Grattafiori et al. (2024). Given that this training paradigm separates the roles of user and assistant, whether a prefix context is provided by the user or generated by the model might influence the model’s recall of edited knowledge. For this analysis, we compare two conditions: (1) prepending a prefix context without any instruction template (original setting) and (2) presenting the context as a user utterance using the user template, followed by the assistant template for generating edited knowledge. We use Llama-3-8B-Instruct and measure the success of knowledge editing based on the appearance of $o^{*}$ and the absence of $o$ within a 10-token window.
@@ 5.5 Effects of Hop Words
Furthermore, as another experiment aimed at examining the impact of hop words, we compared Efficacy between prefix contexts composed solely of hop words (e.g., “Magic Johnson”) and those using fully formed sentences generated from the same hop words (e.g., “Magic Johnson’s impact on the game …”). The results indicated that the difference in Efficacy between these two settings is minimal, averaging around 0.4%, suggesting that hop words alone already substantially contribute to distractiveness. Detailed numerical results can be found in Table 7.
@@ 5.6 Average Contribution Score
We analyze the influence of hop words more quantitatively based on attention scores. Specifically, we define a metric, Average Contribution Score (ACS), as the proportion of prefix contexts in which a hop word receives the highest attention weight among all words in the context, during the last decoding step of knowledge generation
To evaluate ACS, we measure how strongly each token $t_{i}$ in the prefix context attends to the final token $t_{\text{last}}$ by aggregating attention weight across all layers and heads in a pretrained Transformer model.
Let $A_{\ell,h}(t_{i},t_{\text{last}})$ denote the attention weight from token $t_{i}$ to $t_{\text{last}}$ at layer $\ell$ and head $h$. Let $L$ be the number of layers and $H$ be the number of heads per layer. We define the token-level average attention score $\bar{A}_{i\to\text{last}}$ as:
@@
Then the sentence-level informative indicator $I_{n}$ is defined as:
$I_{n}=\begin{cases}1&\text{if}\quad\displaystyle\arg\max_{i\in\text{prefix}}\bar{A}_{i\to\text{last}}=t_{\text{hop}},\\[4.0pt] 0&\text{otherwise}.\end{cases}$ (6)
Finally, we define ACS as the average informativeness over all $N$ samples:
$\mathrm{ACS}=\frac{1}{N}\sum_{n=1}^{N}I_{n}.$ (7)
This value quantifies how often the hop word achieves the highest average attention score, reflecting its informative influence in the prefix context.
In Figure 6, the blue bars compare the ACS of hop words before (darker) and after (lighter) knowledge editing by CoRE. Compared to random chance (red line), hop words receive significantly more attention. However, for $o_{hop}$, which is the most distractive type of hop word, the model pays less attention to them after being edited by CoRE (lighter blue), providing an explanation for CoRE’s context robustness. Conversely, the model pays even greater attention to $o^{*}_{hop}$ after editing. Since $o^{*}_{hop}$ is related to the edited knowledge, this result shows that the edited model learns to extract this useful signal from the context. The red bars in the figure represent the user context setting and show the same pattern. See Appendix E.4 for more details.
@@ E.4 ACS
Recent studies suggest that information flow, particularly the attention from the subject token to the last token of the sentence, plays a crucial role in LLM’s generative performance Geva et al. (2023). Based on this, we further investigated the influence of hop words on knowledge editing performance by measuring the Average Contribution Score (ACS).

### v2

# Context-Robust Knowledge Editing for Language Models

@@ Abstract
Knowledge editing (KE) methods offer an efficient way to modify knowledge in large language models. Current KE evaluations typically assess editing success by considering only the edited knowledge without any preceding contexts. In real-world applications, however, preceding contexts often trigger the retrieval of the original knowledge and undermine the intended edit. To address this issue, we have developed CHED—a benchmark designed to evaluate the context robustness of KE methods. Evaluations on CHED show that they often fail when preceding contexts are present. To mitigate this shortcoming, we introduce CoRE, a KE method designed to strengthen context robustness by minimizing context-sensitive variance in hidden states of the model for edited knowledge. This method not only improves the editing success rate in situations where a preceding context is present but also preserves the overall capabilities of the model. We also provide an in-depth analysis of the differing impacts of preceding contexts when introduced as user utterances versus assistant responses, and we dissect attention-score patterns to assess how specific tokens influence editing success. Our dataset and code are available at https://github.com/holi-lab/CoRE.
Context-Robust Knowledge Editing for Language Models
@@ 2 Related Work
As a first step toward context robust editing, our CoRE method focuses on the locate-then-edit paradigm due to its capacity for supporting mass edits while maintaining model performance. However, we acknowledge the importance of examining and improving context robustness of editing methods in other paradigms and leave this to future work.
@@ 3.2 Contextual Word Selection (Figure 2-B)
We explored additional criteria to identify words that are closely and uniquely associated with the entities in given fact triplets. For instance, we considered hop words with high cosine similarity to the main entity based on BERT embeddings, capturing semantic closeness. Additionally, we measured the probability that a hop word co-occurs with the entity. Table 2 summarizes the criteria considered for hop word selection.
@@
We quantitatively assessed the coherence of the generated prefix contexts using G-Eval Liu et al. (2023) with GPT-4o-mini. The average coherence score across six types of prefix contexts is 3.4 on a scale from 1 to 5, indicating moderate plausibility. Importantly, if we exclude the $o^{*}$ and $o^{*}_{\!hop}$ types, whose content necessarily diverges from the real world facts, the mean coherence rises to 3.8. The lower-than-ideal coherence scores likely result from our selection of primarily low-frequency hop words (as detailed in §3.2), which inherently constrained coherence potential. Nevertheless, because editing success must remain effective regardless of the preceding context’s coherence, we deliberately prioritized distractiveness and consider this level of coherence acceptable. A detailed explanation of the G-Eval process and results is provided in Appendix A.7.
@@ 4.1 Preliminaries
Transformer MLP as a Key-Value Associative Memory
MEMIT interprets MLP layers in Transformers as linear associative memories Anderson (1972); Kohonen (1972), where the weights of the projection layer store key-value associations. For example, when a prompt such as “Tim Cook, who works for” is provided as input, the hidden state of the subject’s last token (i.e., “Cook”) encoded by the first MLP layer serves as the key vector $\mathbf{k}$. As $\mathbf{k}$ passes through the second MLP layer $W_{\text{proj}}$, the stored association relevant to the subject is retrieved and embedded into the output value vector $\mathbf{v}$ that contains information about the associated object (e.g., Apple). At subsequent layers, attention mechanisms refine and propagate this recalled knowledge from the value vector, leading the model to generate the token for $o$ Meng et al. (2022a); Geva et al. (2023).
@@
Key-Value Vector Extraction
A key challenge is constructing the key-value pairs that encode the factual edit $(s,r,o)\to(s,r,o^{*})$. $\mathbf{k}$ and $\mathbf{v}$ are derived from a prompt $p$ that includes $s$ and $r$ and aims to elicit the model’s knowledge. In MEMIT, various prefix contexts $x_{j}$ are prepended to $p$ to improve contextual generalization. Given $N$ prefix contexts, the key vector is derived as $\mathbf{k}=\frac{1}{N}\sum_{j=1}^{N}k(x_{j}+p),$ where $k(\cdot)$ is obtained by extracting the MLP activation at the last subject token from a chosen layer. We defer the full derivation to Appendix B.1.
@@ 5.2 Experimental Settings
In this paper, we focus on comparing locate-then-edit methods, as they reliably handle a large number of edits, including JEEP Shi et al. (2024), EMMET Gupta et al. (2024b), and PMET Li et al. (2023). We also include FT-M Zhang et al. (2024b) as a representative fine-tuning approach. While we experimented with two representative approaches—a meta-learning method MEND Mitchell et al. (2022a) and a weight-preserved method IKE Zheng et al. (2023a)—both achieved only 0–1% edit success under our stricter generation-based metric, effectively amounting to complete editing failure on 1,000 edits. Consequently, we omit them from Table 4. Details about each method, their results, and hyperparameter settings are provided in Appendix F.7.
@@ 5.3 Main Results
Recent language models are typically trained for dialogues with users using instruction templates Touvron et al. (2023); Grattafiori et al. (2024). Given that this training paradigm separates the roles of user and assistant, whether a prefix context is provided by the user or generated by the model might influence the model’s recall of edited knowledge. For this analysis, we compare two conditions: (1) prepending a prefix context without any instruction template (original setting) and (2) presenting the context as a user utterance using the user template, followed by the assistant template for generating edited knowledge. We use Llama-3-8B-Instruct and measure the success of knowledge editing based on the appearance of $o^{*}$ and the absence of $o$ within a 10-token window.
@@ 5.5 Effects of Hop Words
As another experiment for examining the impact of hop words, we compared Efficacy between prefix contexts composed solely of hop words (e.g., “Magic Johnson”) and those using fully formed sentences generated from the same hop words (e.g., “Magic Johnson’s impact on the game …”). The results indicated that the difference in Efficacy between these two settings is minimal, averaging around 0.4%, suggesting that hop words alone already substantially contribute to distractiveness. Detailed numerical results can be found in Table 7.
@@ 5.6 Average Contribution Score
We analyze the influence of hop words more quantitatively based on attention scores. Specifically, we define a metric, Average Contribution Score (ACS), as the proportion of prefix contexts in which a hop word receives the highest attention weight among all words in the context, during the last decoding step of knowledge generation.
More specifically, we measure how strongly the final token $t_{\text{last}}$ attends to each token $t_{i}$ in the prefix context by aggregating $t_{i}$’s attention weights across all layers and heads in a pretrained Transformer model. Formally, let $A_{\ell,h}(t_{i},t_{\text{last}})$ denote the attention weight of token $t_{i}$ received from $t_{\text{last}}$ at layer $\ell$ and head $h$. Let $L$ be the number of layers and $H$ be the number of heads per layer. We define the token-level average attention score $\bar{A}_{i\to\text{last}}$ as:
@@
Given a knowledge editing case with a prefix context containing a hop word $t_{hop}$, the indicator $I$ of whether the hop word receives the highest attention is defined as:
$I=\begin{cases}1&\text{if }\operatorname*{argmax}_{i\in\text{prefix}}\bar{A}_{i\to\text{last}}=t_{\text{hop}},\\[4.0pt] 0&\text{otherwise}.\end{cases}$ (6)
Finally, we define ACS as the percentage of test cases where hop words receive the highest attention:
$\mathrm{ACS}=\frac{1}{N}\sum_{n=1}^{N}I_{n},$ (7)
where $N$ is the number of test cases. This value quantifies reflects the degree of influence of hop words on knowledge recall.
In Figure 6, the blue bars compare the ACS of hop words before (darker) and after (lighter) knowledge editing by CoRE. Compared to random chance (red line), hop words receive significantly more attention. However, for $o_{hop}$, which is the most distractive type of hop words, the model pays less attention to them after being edited by CoRE (lighter blue), explaining CoRE’s context robustness. Conversely, the model pays even greater attention to $o^{*}_{hop}$ after editing. Since $o^{*}_{hop}$ is related to the edited knowledge, it provides a useful signal for edited knowledge. The results show that CoRE does not simply reduce the model’s attention to prefix contexts; rather, it improves the model’s ability to attend less to distractive information and more to useful information in the context. The red bars in the figure represent the user context setting and show the same pattern. See Appendix E.4 for more details.
@@ E.4 ACS
Recent studies suggest that information flow, particularly the attention from the subject token to the last token of the sentence, plays a crucial role in LLM’s generative performance Geva et al. (2023). Based on this, we further investigated the influence of hop words on knowledge editing performance by measuring the Average Contribution Score (ACS). If a hop word spans multiple tokens, we compute its total impact by summing the contributions of each constituent token.